# When Person Re-Identification Meets Event Camera: A Benchmark Dataset and An Attribute-guided Re-Identification Framework

### v1

@@ 4.3 Input Representation
where $\mathcal{V}(\cdot)$ denotes the CLIP visual encoder. To incorporate identity-level textual guidance, we construct an ID-specific prompt: “A photo of a $[X]_{1},[X]_{2},\cdots,[X]_{\text{Num}}$ person”, where each $[X]_{y_{i}}$ ($y_{i}\in\{1,\dots,\text{Num}\}$) is a learnable token with the same dimension as the word embedding and is independently learned for each identity. The textual prompt is fed into the CLIP text encoder $\mathcal{T}(\cdot)$ to obtain the identity-specific text feature:
@@ 4.4 Attribute Guided Prompt Learning
$\bullet$ Positive-Negative Attribute Prompts. To further enhance the discriminative capacity of identity representations with fine-grained semantic cues, we propose a Positive-Negative Attribute Prompting (PNAP) mechanism. This design is inspired by the learnable context prompt strategy in [59, 21], and is further motivated by [51], which demonstrates the benefits of incorporating attribute-level semantic priors for multi-modal object re-identification. Specifically, we employ a pre-trained pedestrian attribute recognition model [48], denoted as $\mathcal{PAR}(\cdot)$, to predict attribute labels for a given input sequence. Based on the predicted results, we construct a positive attribute prompt $A^{\text{pos}}$, which contains all the predicted attributes (e.g., “Male, Jacket, Bald”), and a negative attribute prompt $A^{\text{neg}}$ by negating the unpredicted attributes (e.g., “Not Female, Not Short Sleeves, Not Long Hair”). Both types of prompts are encoded into textual embeddings using the frozen CLIP text encoder $\mathcal{T}(\cdot)$. Following the prompt tuning paradigm of CoOp [59], we further introduce a set of learnable context vectors: $P^{\text{pos}}$ for positive prompts and $P^{\text{neg}}$ for negative ones. These are used to adapt the attribute prompts to the target person ReID task. The enriched prompts are then passed through a fully connected (FC) layer and concatenated with the visual features $\hat{F}_{R}$ and $\hat{F}_{E}$ to yield prompt-enhanced identity representations $F_{R}^{0}$ and $F_{E}^{0}$. The overall process can be summarized as:
@@
$\bullet$ Cross Modal Prompts. Inspired by the Visual Prompt Tuning paradigm [20], we design a Cross Modal Prompting (CMP) mechanism to facilitate bidirectional interaction between RGB and Event modalities at the feature level, which introduces a set of learnable prompts to enable early-stage alignment across modalities. Concretely, we initialize a set of modality-specific prompt tokens $CMP^{0}_{R}$ for the RGB branch and project them into the Event branch via an FC layer to obtain $CMP^{0}_{E}$. These prompts are concatenated with the visual features $F_{R}^{0}$ and $F_{E}^{0}$, resulting in the prompt-injected features $F_{R}^{1}$ and $F_{E}^{1}$, respectively:
@@ 4.5 Training Strategy
$\bullet$ Stage 1: Text Prompt Alignment. In the first stage, we follow the paradigm of CLIP-ReID and introduce ID-specific learnable prompts to model ambiguous identity-related semantics. For each identity, we learn a set of $n$ text tokens ${[X]_{1},[X]_{2},\cdots,[X]_{n}}$, which are concatenated to form a textual description in the format of “A photo of a $[X]_{1},[X]_{2},\cdots,[X]_{n}$ person”. These prompts are optimized while keeping both the visual encoder $\mathcal{V}(\cdot)$ and text encoder $\mathcal{T}(\cdot)$ frozen. The optimization is guided by two symmetric contrastive losses, vision-to-text ($L{v2t}$) and text-to-vision ($L{t2v}$), which encourage cross-modal alignment between video features and the learnable textual embeddings. Formally, given a mini-batch of size $|B|$, the overall objective is defined as:
@@
1 Stage 1: Text Prompt Alignment;
@@ 5       Compute contrastive loss $L_{v2t}+L_{t2v}$;
6
7 Stage 2: Multimodal Prompt Alignment;
@@ 12       Compute $L_{v2t}+L_{t2v}$;
13
14 Stage 3: Visual-Modal Fine-Tuning with Attribute Prompts;
@@ 5 Experiments
Methods Modality Publish Code mAP Rank-1 Rank-5 Rank-10
#01 OSNet [58] V+E ICCV2019 URL 23.7 49.1 65.4 72.3
#02 TCLNet [16] V+E ECCV2020 URL 55.8 77.4 89.0 93.1
#03 AP3D [14] V+E ECCV2020 URL 66.9 86.5 95.6 96.5
#04 MGH [53] V+E CVPR2020 URL 43.2 70.9 89.4 92.7
#05 STMN [11] V+E ICCV2021 URL 42.1 73.8 - -
#06 PSTA [49] V+E ICCV2021 URL 68.2 82.3 90.8 94.5
#07 GRL [29] V+E CVPR2021 URL 38.9 62.6 78.0 83.6
#08 BiCnet-TKS [17] V+E CVPR2021 URL 50.8 80.5 89.62 92.45
#09 SINet [6] V+E CVPR2022 URL 50.2 77.4 92.8 96.2
#10 DCCT [30] V+E TNNLS2023 URL 24.6 42.7 64.9 75.5
#11 CLIP-ReID [25] V+E AAAI2023 URL 49.2 73.0 85.5 91.5
#12 SDCL [9] V+E CVPR2023 URL 54.2 69.3 83.8 87.1
#13 TF-CLIP [54] V+E AAAI2024 URL 56.9 78.6 91.8 94.3
#14 DeMo [50] V+E AAAI2025 URL 59.4 75.7 90.1 92.8
#15 CLIMB-ReID [55] V+E AAAI2025 URL 68.3 85.2 92.8 95.8
#16 TriPro-ReID (Ours) V+E - URL 69.3 88.6 94.3 95.4
@@
#01 GRL [29] V+E 82.8 88.7
#02 OSNet [58] V+E 81.9 87.7
#03 SRS-Net [41] V+E 83.8 89.3
#04 STMN [11] V+E 83.4 89.0
#05 CTL [28] V+E 85.3 89.6
#06 PSTA [49] V+E 85.1 89.9
#07 SDCL [9] V+E 86.5 91.1
@@ 5.1 Datasets and Evaluation Metric
To evaluate the performance, we conduct a comprehensive benchmark of 10 pedestrian attribute re-identification methods, representing the most important models in the field of pedestrian attribute re-identification. Since there is no available RGB-Event person ReID dataset, Cao et al. [9] generate events from MARS [57]. We adopt this simulated dual-modal MARS RGB-Event dataset (MARS) and our proposed EvReID dataset for evaluation. Following previous works, we adopt the mean Average Precision (mAP) and Cumulative Matching Characteristics (CMC) at Rank-$\mathcal{K}$ ($\mathcal{K}=1,5,10$) as our evaluation metrics.
@@ 5.2 Implementation Details
We use CLIP-ReID [26] as our baseline method, and the two modalities adopt independent backbone parameters. Our method consists of three training stages. In the first stage, we only optimize the ID-related context prompts. In the second stage, we introduce 20 randomly initialized cross-modal visual prompts into the RGB branch and generate Event-modality visual prompts via a projection layer. During this stage, only the cross-modal visual prompts and the projection layer are optimized. In the third stage, the text encoder and text context prompts remain frozen. We introduce learnable attribute context prompts and PNAP, and optimize them along with CMP and the visual backbone. SGD is used as the optimizer in the first two stages, while AdamW [31] is employed in the final stage. The learning rates are set to $3.5\times 10^{-3}$, $3.5\times 10^{-3}$, and $65e-$ 6, respectively. All training is conducted on an NVIDIA RTX 3090 GPU, with batch sizes of 64, 64, and 8, respectively. More details can be found in our source code.
@@ 6 Conclusion and Future Works
This paper presents TriPro-ReID, a novel approach for RGB-Event-based person re-identification, which integrates an attribute-guided framework. Our method demonstrates superior performance, addressing key challenges in existing dual-modality person ReID. Furthermore, we introduce a new dataset, EvReID, designed to foster progress in this domain by providing a comprehensive benchmark that spans diverse seasons, scenes, and lighting conditions. In addition, we retrain 15 state-of-the-art methods on this dataset, contributing to the ongoing development of RGB-Event-based person re-identification. In the future, we plan to expand the scale of the dataset and investigate the impact of temporal dynamics on model performance. Additionally, the use of more advanced large foundation models [45] will be explored to enhance the overall accuracy.

### v2

@@ 4.3 Input Representation
where $\mathcal{V}(\cdot)$ denotes the CLIP visual encoder. To incorporate identity-level textual guidance, we construct an ID-specific prompt: “A photo of a $X]_{1},[X]_{2},\cdots,[X]_{\text{Num}}$ person”, where each $X]_{y_{i}}$ ($y_{i}\in\{1,\dots,\text{Num}\}$) is a learnable token with the same dimension as the word embedding and is independently learned for each identity. The textual prompt is fed into the CLIP text encoder $\mathcal{T}(\cdot)$ to obtain the identity-specific text feature:
@@ 4.4 Attribute Guided Prompt Learning
$\bullet$ Positive-Negative Attribute Prompts. To further enhance the discriminative capacity of identity representations with fine-grained semantic cues, we propose a Positive-Negative Attribute Prompting (PNAP) mechanism. This design is inspired by the learnable context prompt strategy in [zhou2022learning, khattak2023maple], and is further motivated by [wang2025idea], which demonstrates the benefits of incorporating attribute-level semantic priors for multi-modal object re-identification. Specifically, we employ a pre-trained pedestrian attribute recognition model [wang2024spatio], denoted as $\mathcal{PAR}(\cdot)$, to predict attribute labels for a given input sequence. Based on the predicted results, we construct a positive attribute prompt $A^{\text{pos}}$, which contains all the predicted attributes (e.g., “Male, Jacket, Bald”), and a negative attribute prompt $A^{\text{neg}}$ by negating the unpredicted attributes (e.g., “Not Female, Not Short Sleeves, Not Long Hair”). Both types of prompts are encoded into textual embeddings using the frozen CLIP text encoder $\mathcal{T}(\cdot)$. Following the prompt tuning paradigm of CoOp [zhou2022learning], we further introduce a set of learnable context vectors: $P^{\text{pos}}$ for positive prompts and $P^{\text{neg}}$ for negative ones. These are used to adapt the attribute prompts to the target person ReID task. The enriched prompts are then passed through a fully connected (FC) layer and concatenated with the visual features $\hat{F}_{R}$ and $\hat{F}_{E}$ to yield prompt-enhanced identity representations $F_{R}^{0}$ and $F_{E}^{0}$. The overall process can be summarized as:
@@
$\bullet$ Cross Modal Prompts. Inspired by the Visual Prompt Tuning paradigm [jia2022visual], we design a Cross Modal Prompting (CMP) mechanism to facilitate bidirectional interaction between RGB and Event modalities at the feature level, which introduces a set of learnable prompts to enable early-stage alignment across modalities. Concretely, we initialize a set of modality-specific prompt tokens $CMP^{0}_{R}$ for the RGB branch and project them into the Event branch via an FC layer to obtain $CMP^{0}_{E}$. These prompts are concatenated with the visual features $F_{R}^{0}$ and $F_{E}^{0}$, resulting in the prompt-injected features $F_{R}^{1}$ and $F_{E}^{1}$, respectively:
@@ 4.5 Training Strategy
$\bullet$ Stage 1: Text Prompt Alignment. In the first stage, we follow the paradigm of CLIP-ReID and introduce ID-specific learnable prompts to model ambiguous identity-related semantics. For each identity, we learn a set of $n$ text tokens $X]_{1},[X]_{2},\cdots,[X]_{n}$}, which are concatenated to form a textual description in the format of “A photo of a $X]_{1},[X]_{2},\cdots,[X]_{n}$ person”. These prompts are optimized while keeping both the visual encoder $\mathcal{V}(\cdot)$ and text encoder $\mathcal{T}(\cdot)$ frozen. The optimization is guided by two symmetric contrastive losses, vision-to-text (Lv2tL{v2t}) and text-to-vision (Lt2vL{t2v}), which encourage cross-modal alignment between video features and the learnable textual embeddings. Formally, given a mini-batch of size $|B|$, the overall objective is defined as:
@@
1
20.5em Stage 1: Text Prompt Alignment;
@@ 6   Compute contrastive loss $L_{v2t}+L_{t2v}$;
7
8
90.5em Stage 2: Multimodal Prompt Alignment;
@@ 14   Compute $L_{v2t}+L_{t2v}$;
15
16
170.5em Stage 3: Visual-Modal Fine-Tuning with Attribute Prompts;
@@ 5 Experiments
Methods Modality Publish mAP Rank-1 Rank-5 Rank-10
#01 AP3D [gu2020AP3D] V ECCV2020 65.4 83.0 92.1 95.3
#02 PSTA [PSTA] V ICCV2021 63.3 85.2 92.5 95.3
#03 GRL [liu2021watching] V CVPR2021 34.5 58.2 73.6 84.0
#04 SINet [bai2022SINet] V CVPR2022 62.7 83.0 92.4 96.5
#05 CLIP-ReID [li2022clip] V AAAI2023 49.9 68.8 81.0 84.0
#01 AP3D [gu2020AP3D] E ECCV2020 40.6 67.3 80.8 86.2
#02 PSTA [PSTA] E ICCV2021 37.1 61.0 77.4 84.6
#03 GRL [liu2021watching] E CVPR2021 38.9 62.6 78.0 83.6
#04 SINet [bai2022SINet] E CVPR2022 40.1 67.0 81.5 85.2
#05 CLIP-ReID [li2022clip] E AAAI2023 30.4 52.5 70.3 79.1
#01 OSNet [zhou2019omniscalefeaturelearningperson] V+E ICCV2019 23.7 49.1 65.4 72.3
#02 TCLNet [TCLNet] V+E ECCV2020 55.8 77.4 89.0 93.1
#03 AP3D [gu2020AP3D] V+E ECCV2020 66.9 86.5 95.6 96.5
#04 MGH [DBLP:conf/cvpr/YanQC0ZT020] V+E CVPR2020 43.2 70.9 89.4 92.7
#05 STMN [eom2021video] V+E ICCV2021 42.1 73.8 - -
#06 PSTA [PSTA] V+E ICCV2021 68.2 82.3 90.8 94.5
#07 GRL [liu2021watching] V+E CVPR2021 38.9 62.6 78.0 83.6
#08 BiCnet-TKS [BiCnet-TKS] V+E CVPR2021 50.8 80.5 89.6 92.5
#09 SINet [bai2022SINet] V+E CVPR2022 67.1 83.4 93.0 95.6
#10 DCCT [liu2023deeply] V+E TNNLS2023 24.6 42.7 64.9 75.5
#11 CLIP-ReID [li2022clip] V+E AAAI2023 49.2 73.0 85.5 91.5
#12 SDCL [cao2023event] V+E CVPR2023 54.2 69.3 83.8 87.1
#13 TF-CLIP [tfclip] V+E AAAI2024 56.9 78.6 91.8 94.3
#14 DeMo [wang2025decoupled] V+E AAAI2025 59.4 75.7 90.1 92.8
#15 CLIMB-ReID [yu2025climb] V+E AAAI2025 68.3 85.2 92.8 95.8
#16 TriPro-ReID (Ours) V+E - 69.3 88.6 94.3 95.4
@@
#01 OSNet [zhou2019omniscalefeaturelearningperson] V 81.4 87.3
#02 GRL [liu2021watching] V 84.8 91.0
#03 STMN [eom2021video] V 84.5 90.5
#04 PSTA [PSTA] V 85.8 91.5
#05 TF-CLIP [tfclip] V 89.4 93.0
#06 CLIMB-ReID [yu2025climb] V 89.7 93.3
#01 TCLNet [TCLNet] E 38.2 25.3
#02 GRL [liu2021watching] E 27.7 16.7
#03 BiCnet-TKS [BiCnet-TKS] E 30.9 17.3
#04 STMN [eom2021video] E 22.4 10.0
#01 OSNet [zhou2019omniscalefeaturelearningperson] V+E 81.9 87.7
#02 GRL [liu2021watching] V+E 82.8 88.7
#03 SRS-Net [wang2020simple] V+E 83.8 89.3
#04 STMN [eom2021video] V+E 83.4 89.0
#05 PSTA [PSTA] V+E 85.1 89.9
#06 SDCL [cao2023event] V+E 86.5 91.1
@@ 5.1 Datasets and Evaluation Metric
To evaluate the performance, we conduct a comprehensive benchmark of 10 pedestrian attribute re-identification methods, representing the most important models in the field of pedestrian attribute re-identification. Since there is no available RGB-Event person ReID dataset, Cao et al. [cao2023event] generate events from MARS [zheng2016mars]. We adopt this simulated dual-modal MARS RGB-Event dataset (MARS) and our proposed EvReID dataset for evaluation. Following previous works, we adopt the mean Average Precision (mAP) and Cumulative Matching Characteristics (CMC) at Rank-$\mathcal{K}$ ($\mathcal{K}=1,5,10$) as our evaluation metrics.
@@ 5.2 Implementation Details
We use CLIP-ReID [li2023clip] as our baseline method, and the two modalities adopt independent backbone parameters. Our method consists of three training stages. In the first stage, we only optimize the ID-related context prompts. In the second stage, we introduce 20 randomly initialized cross-modal visual prompts into the RGB branch and generate Event-modality visual prompts via a projection layer. During this stage, only the cross-modal visual prompts and the projection layer are optimized. In the third stage, the text encoder and text context prompts remain frozen. We introduce learnable attribute context prompts and PNAP, and optimize them along with CMP and the visual backbone. SGD is used as the optimizer in the first two stages, while AdamW [loshchilov2017decoupled] is employed in the final stage. The learning rates are set to $3.5\times 10^{-3}$, $3.5\times 10^{-3}$, and 5e65e-6, respectively. All training is conducted on an NVIDIA RTX 3090 GPU, with batch sizes of 64, 64, and 8, respectively. More details can be found in our source code.
@@ 6 Conclusion and Future Works
This paper presents TriPro-ReID, a novel approach for RGB-Event-based person re-identification, which integrates an attribute-guided framework. Our method demonstrates superior performance, addressing key challenges in existing dual-modality person ReID. Furthermore, we introduce a new dataset, EvReID, designed to foster progress in this domain by providing a comprehensive benchmark that spans diverse seasons, scenes, and lighting conditions. In addition, we retrain 15 state-of-the-art methods on this dataset, contributing to the ongoing development of RGB-Event-based person re-identification. In the future, we plan to expand the scale of the dataset and investigate the impact of temporal dynamics on model performance. Additionally, the use of more advanced large foundation models [wang2023MMPTMSurvey] will be explored to enhance the overall accuracy.